# {textblock*}

### v1

@@ Abstract
Within the Electronic Design Automation (EDA) domain, AI-driven solutions have emerged as formidable tools, yet they typically augment rather than redefine existing methodologies. These solutions often repurpose deep learning models from other domains—such as vision, text, and graph analytics—applying them to circuit design without tailoring to the unique complexities of electronic circuits. Such an “AI4EDA” approach falls short of achieving a holistic design synthesis and understanding, overlooking the intricate interplay of electrical, logical, and physical facets of circuit data. This perspective paper argues for a paradigm shift from AI4EDA towards AI-native EDA, integrating AI at the core of the design process. Pivotal to this vision is the development of a multimodal circuit representation learning technique, poised to provide a comprehensive understanding by harmonizing and extracting insights from varied data sources, such as functional specifications, RTL designs, circuit netlists, and physical layouts.
We champion the creation of large circuit models (LCMs) that are inherently multimodal, crafted to decode and express the rich semantics and structures of circuit data, thus fostering more resilient, efficient, and inventive design methodologies. Embracing this AI-native philosophy, we foresee a trajectory that transcends the current innovation plateau in EDA, igniting a profound “shift-left” in electronic design methodology. The envisioned advancements herald not just an evolution of existing EDA tools but a revolution, giving rise to novel instruments of design—tools that promise to radically enhance design productivity and inaugurate a new epoch where the optimization of circuit performance, power, and area (PPA) is achieved not incrementally, but through leaps that redefine the benchmarks of electronic systems’ capabilities.
@@ 1.1 The Rise of Foundation Models
Large foundation models represent a significant leap in AI. These models, typically pre-trained on web-scale datasets using self-supervision techniques [1], have been adapted to excel in a wide array of downstream tasks. In domains rich with structured data, such as natural language processing (NLP) and computer vision (CV), these models have not only set new benchmarks but have fundamentally redefined the realms of possibility.
In NLP, models like BERT [2] and its derivatives, including RoBERTa [3] and T5 [4], have revolutionized language understanding, especially in contextual interpretation of text, thereby enhancing complex language-based tasks. Concurrently, the decoder-only GPT series [5] has shown remarkable versatility, excelling in diverse tasks from creative writing to code generation and pointing towards the burgeoning potential of artificial general intelligence (AGI). In visual AI, self-supervised foundation models [6, 7, 8] have achieved competitive performances in image understanding tasks, rivaling fully supervised approaches.
@@ 1.3 The Feasibility and Promises of AI-Native LCMs
Moreover, the landscape of circuit types, though vast, is marked by commonalities that transcend individual designs. Processors, domain accelerators such as digital signal processors (DSPs) and AI accelerators, communication modules, and other core components display a pattern of design module reuse—arithmetic units, various decoders, and cryptographic cores are just a few examples. This consistency provides a predictable pattern—akin to an inductive bias—that is conducive to the application of machine learning models.
@@ 2.1 Core Objectives and Complexities in EDA
In the fiercely competitive realm of electronic product development, reducing time-to-market (TTM) is paramount. The rapid evolution of consumer electronics, exemplified by the yearly refresh cycles of smartphones and wearables, underscores the urgency to expedite product launches to capture market share and meet consumer expectations. This pressure significantly impacts the EDA process, where the need for TTM can sometimes compromise design thoroughness, leading to potential flaws. For instance, under the gun to release the next generation of microprocessors, teams may bypass exhaustive verification in favor of meeting launch windows, risking the introduction of bugs into the final product. When such issues are not amendable through engineering change orders (ECO) [25], they necessitate a costly and time-consuming redesign, further exacerbating time-to-market pressures. This cycle highlights the crucial need for EDA solutions that not only streamline design and verification processes but also ensure design accuracy from the outset.
@@ 2.2 EDA for Front-End Design
Figure 2: A typical front-end design flow.
A typical front-end design flow, also known as logic design, is shown in Figure 2, in which the design specification is transformed into a logic netlist. The front-end design flow begins with a design specification, followed by architecture exploration. Subsequently, HDLs are created to translate the design into a form suitable for implementation, typically at the RTL abstraction level. The introduction of hardware construction languages (like Chisel [27]) and high-level synthesis (HLS) adds a new dimension to front-end design and offers more flexibility and efficiency in addressing the complexities of modern front-end design.
After RTLs are created or generated from HLS tools, designers first use static analysis tools such as lint to identify potential errors and then apply various verification techniques, including logic simulation, emulation, and various formal methods (e.g., model checking). These techniques collectively contribute to validating the functionalities of the RTL design faithfully following the design specification. The verification and testing processes, spanning various transformations and stages, are integral components of design flows, typically consuming between 60% to 70% of the total engineering efforts allocated. This substantial investment underscores their critical role in ensuring the functionality and reliability of circuit designs. Across diverse abstracts of circuit designs, a plethora of verification techniques are employed, reflecting the nuanced requirements and challenges encountered at each stage of development. One notable approach involves C-RTL equivalence checking, where RTL implementations are rigorously compared against C-based specification models. This verification method, as evidenced by studies such as [28, 29], is frequently applied, particularly in the context of data-path intensive designs, as highlighted by Hector from Synopsis[30]. Given that circuit designs within this abstract primarily encapsulate hardware behavior while abstracting concrete physical details, theorem provers and SMT solvers emerge as pivotal tools for enhancing verification efficacy [31, 32].
@@ 2.3 EDA for Back-End Design
Figure 3: A typical back-end design flow.
For modern chip design, the back-end design flow, also referred to as layout design, is depicted in Figure 3, transitioning from a gate-level or generic technology (GTech) netlist to a finalized layout [37].
@@ 2.4.1 EDA for Standard Cells
Standard cells, the building blocks of digital ICs, follow predefined structures that align with a library’s specifications, enabling their reuse across diverse designs. The focus of EDA tools in standard cell design is primarily on automating the layout generation process, encompassing crucial steps like placement and in-cell routing, as shown in Figure 4.
The placement process is dedicated to determining the optimal transistor locations within a cell to maximize space utilization while maintaining functionality and performance integrity. Innovations in placement strategies, as highlighted in [39, 40], have introduced methods to expedite this intricate procedure while ensuring routability and design efficiency. In contrast, in-cell routing tackles the intricate task of establishing connections within the cell, a process complicated by the rigorous area constraints of standard cells. This stage demands specialized routing solutions, distinct from those applied to broader digital circuits, to navigate the tight confines of cell layouts. Contributions from [41, 42] have provided targeted approaches to in-cell routing, addressing the unique challenges of standard cell design.
@@ 2.4.2 EDA for Datapath Circuits
Figure 5: A typical datapath circuits design flow.
@@
These tools adopt a comprehensive strategy for refining datapath circuits, ensuring seamless integration and peak efficiency among components. As depicted in Figure 5, the design journey initiates with pinpointing a target application and its corresponding architectural design, thereby defining a broad and intricate design space. This space might include, for example, CPU tasks like SPEC2017 benchmarks or GPU tasks such as matrix multiplication, targeting either CPU or GPU architectures accordingly. The design space diverges into two principal domains: the application space, outlining application-specific parameters like dataflow patterns or neural network mapping strategies, and the architecture space, detailing the structural and resource parameters, such as CPU issue width or the quantity of MACs in a neural processing unit (NPU).
The intersection of parameters from these domains establishes a ’design point’, which, upon post-compilation application mapping, is subjected to thorough evaluation and validation via cutting-edge EDA tools. This rigorous process iteratively explores and assesses new design points until the exploration goals are achieved.
@@ 2.4.3 EDA for Analog Circuits
Figure 6 illustrates a typical analog IC design flow, starting with a detailed set of circuit specifications covering area, power, and performance requirements. The front-end design phase is crucial, establishing the pre-layout circuit netlist that defines the circuit’s functionalities via meticulous topology design and device sizing. This phase sets the foundation for the circuit’s operational features and optimization criteria.
@@ 3.1.1 Pre-RTL ML Methods
In the realm of tensor applications, HASCO [77] employs ML for DSE. This methodology optimizes both software programs and hardware accelerators, showcasing ML’s capacity to bridge the gap between software and hardware domains to achieve optimized system performance.
@@ 3.1.2 RTL-Stage ML Methods
Power modeling at the RTL stage has also attracted lots of attention. There are two primary categories: design-time power estimation and runtime on-chip power modeling. For design-time estimation, PRIMAL [84] stands out for offering per-cycle power evaluations tailored to each target design, alongside other notable ML-based approaches [85, 86]. For runtime power modeling, DEEP [87] introduces an efficient on-chip model that incorporates low-overhead hardware design, utilizing ML to identify power-correlated RTL signals, or ’power proxies.’ This method, along with other ML-based on-chip power modeling solutions like [88, 89], demonstrates the potential of ML in creating dynamic power models that adapt to real-time conditions. Moreover, APOLLO [90] presents a versatile solution applicable to both design-time and runtime scenarios. Simmani [91] and the early power modeling work [92] focus on fast power emulation on FPGA and other platforms, highlighting the broader applicability of ML methods in facilitating efficient power analysis during the design phase.
@@ 3.1.4 Layout-Stage ML Methods
Placement Stage Enhancements with ML: The placement stage, which determines the optimal locations of macros and standard cells in the layout, is pivotal for achieving desired design metrics. Early applications of ML aimed to augment traditional placement strategies. PADE [110] incorporates support vector machines (SVM) and neural networks for datapath extraction and evaluation, facilitating datapath-aware placement strategies. DREAMPlace, developed by Lin et al. [111], conceptualizes the placement challenge as akin to training a neural network, thus accelerating the global placement process by harnessing GPU computing capabilities. Building on DREAMPlace, Agnesina et al. [112] apply multi-objective Bayesian optimization for macro placement design space exploration, demonstrating the potential of ML in enhancing macro-placement outcomes.
@@
Sign-Off Enhancements with ML: During the routing and sign-off stages, the precision of sign-off timing, especially using the path-based static timing analysis (PBA), becomes crucial. However, the PBA process is time-consuming, leading to the application of machine learning models for predicting path-based timing based on quicker graph-based analysis (GBA) results. The pioneering work by Kahng et al. [128] was instrumental in predicting PBA from GBA using carefully engineered features and a tree-based model. Subsequent studies, such as  [123, 129], have delved into various machine learning models, including transformers and GNN, to enhance the accuracy of GBA-PBA predictions.
@@
Manufacturability Enhancements with ML: In the field of design for manufacturing (DFM), leveraging ML has become pivotal for bolstering the reliability of lithography and manufacturing processes, with layout patterns often analyzed as images. Studies like GAN-SRAF [133], GAN-OPC [134], Develset [135], and L2O-ILT [136] use various ML methods to improve mask synthesis printability. Other works, such as those by Watanabe et al. [137], Ye et al. [138], Lin et al. [139] and Chen et al. [140], focus on lithography modeling to simulate printed patterns from mask clips. For identifying layout patterns prone to printing failures like shorts or opens, ML-enhanced lithography hotspot detection is explored in various studies. For example, Yang et al. [141] propose to extract layout features with discrete cosine transform and utilize a CNN architecture for hotspot detection. The performance is further improved with the proposed bias learning algorithm because of the imbalanced dataset. Inspired by the object detection problem in computer vision, Chen et al. [142] propose to detect multiple hotspots within large layouts simultaneously. In [143], the binarized neural network is utilized to speed up the hotspot detection flow. New network architecture is designed based on residual networks to achieve higher detection accuracy and performance. Additionally, ML further contributes to yield estimation and analysis, as seen in works like Ciccazzo et al. [144], Nakata et al. [145], and Alawieh et al. [146].
@@ 3.2 Reinforcement Learning in EDA
Reinforcement learning (RL) in EDA has emerged as a powerful method for navigating the expansive solution spaces inherent in logic synthesis and physical design, often uncovering innovative solutions that surpass traditional, intuition-based approaches. Innovations like Synopsys.ai underscore this trend, showcasing AI-driven methodologies that enhance PPA metrics across the design spectrum.
@@ 3.3 Leveraging Large Language Models in EDA
This section delves into the use of LLMs for RTL code generation—a key area of focus. It categorizes the research into benchmarking efforts, the use of commercial LLMs, and the development of specialized open-source LLMs through fine-tuning. Additionally, we explore LLM applications in generating hardware security codes and EDA scripts.
@@ 3.3.2 Enhancing Verification with LLMs
Functional Verification through LLMs: LLMs have made significant strides in functional verification by translating natural language specifications into SystemVerilog assertions (SVAs). This process ensures that RTL implementations adhere to their intended specifications. Notably, [174, 175] leverage human-written specification sentences alongside RTL designs to generate precise SVAs. AssertLLM [176] takes a proactive approach by generating assertions directly from comprehensive specification documents, even before the RTL design phase. This method is complemented by a benchmark set that pairs natural language specifications with golden RTL implementations, offering a robust framework for evaluating assertion generation.

### v2

@@ Abstract
Within the Electronic Design Automation (EDA) domain, AI-driven solutions have emerged as formidable tools, yet they typically augment rather than redefine existing methodologies. These solutions often repurpose deep learning models from other domains such as vision, text, and graph analytics applying them to circuit design without tailoring to the unique complexities of electronic circuits. Such an “AI4EDA” approach falls short of achieving a holistic design synthesis and understanding, overlooking the intricate interplay of electrical, logical, and physical facets of circuit data. This paper argues for a paradigm shift from AI4EDA towards AI-native EDA, integrating AI at the core of the design process. Pivotal to this vision is the development of a multimodal circuit representation learning technique, poised to provide a comprehensive understanding by harmonizing and extracting insights from varied data sources, such as functional specifications, RTL designs, circuit netlists, and physical layouts.
We champion the creation of large circuit models (LCMs) that are inherently multimodal, crafted to decode and express the rich semantics and structures of circuit data, thus fostering more resilient, efficient, and inventive design methodologies. Embracing this AI-native philosophy, we foresee a trajectory that transcends the current innovation plateau in EDA, igniting a profound “shift-left” in electronic design methodology. The envisioned advancements herald not just an evolution of existing EDA tools but a revolution, giving rise to novel instruments of design tools that promise to radically enhance design productivity and inaugurate a new epoch where the optimization of circuit performance, power, and area (PPA) is achieved not incrementally, but through leaps that redefine the benchmarks of electronic systems’ capabilities.
@@ 1.1 The Rise of Foundation Models
Large foundation models represent a significant leap in AI. These models, typically pre-trained on web-scale datasets using self-supervision techniques [1], have been adapted to excel in a wide array of downstream tasks. In the fields of natural language processing (NLP) and computer vision (CV), these models have not only set new benchmarks but have fundamentally redefined the realms of possibility.
In NLP, models like BERT [2] and its derivatives, including RoBERTa [3] and T5 [4], have revolutionized language understanding, especially in contextual interpretation of text, thereby enhancing complex language-based tasks. Concurrently, the decoder-only GPT series [5] has shown remarkable versatility, excelling in diverse tasks from creative writing to code generation and pointing towards the burgeoning potential of artificial general intelligence (AGI). In CV area, self-supervised foundation models [6, 7, 8] have achieved competitive performances in image understanding tasks, rivaling fully supervised approaches.
@@ 1.3 The Feasibility and Promises of AI-Native LCMs
Moreover, the landscape of circuit types, though vast, is marked by commonalities that transcend individual designs. Processors, domain accelerators (e.g. digital signal processors (DSPs) and AI accelerators), communication modules, and other core components display a pattern of design module reuse. Examples of these reusable modules include arithmetic units, various decoders, and cryptographic cores. This consistency provides a predictable pattern—akin to an inductive bias—that is conducive to the application of machine learning models.
@@ 2.1 Core Objectives and Complexities in EDA
In the fiercely competitive realm of electronic product development, reducing time-to-market (TTM) is paramount. The rapid evolution of consumer electronics, exemplified by the yearly refresh cycles of smartphones and wearables, underscores the urgency to expedite product launches to capture market share and meet consumer expectations. This pressure significantly impacts the EDA process, where the need for TTM can sometimes compromise design thoroughness, leading to potential flaws. For instance, under the gun to release the next generation of microprocessors, teams may bypass exhaustive verification in favor of meeting launch windows, risking the introduction of bugs into the final product. When such issues are not amendable through engineering change orders (ECO) [25], they necessitate a costly and time-consuming redesign, further exacerbating time-to-market pressures. Therefore, this cycle highlights the crucial need for EDA solutions that not only streamline design and verification processes but also ensure design accuracy from the outset.
@@ 2.2 EDA for Front-End Design
Figure 2: Typical front-end and back-end design flows.
A typical front-end design flow, also known as logic design, is shown in Fig. 2 a), in which the design specification is transformed into a logic netlist. The front-end design flow begins with a design specification, followed by architecture exploration. Subsequently, HDLs are created to translate the design into a form suitable for implementation, typically at the RTL abstraction level. The introduction of hardware construction languages (like Chisel [27]) and C/C++ high-level synthesis (HLS) adds a new dimension to front-end design and offers more flexibility and efficiency in addressing the complexities of modern front-end design.
After RTLs are created or generated from HLS tools, designers first use static analysis tools such as Lint [28] to identify potential errors and then apply various verification techniques, including logic simulation, emulation, and various formal methods (e.g., model checking). These techniques collectively contribute to validating the functionalities of the RTL design faithfully following the design specification. The verification and testing processes, spanning various transformations and stages, are integral components of design flows, typically consuming between 60% to 70% of the total engineering efforts allocated. This substantial investment underscores their critical role in ensuring the functionality and reliability of circuit designs. Across diverse abstracts of circuit designs, a plethora of verification techniques are employed, reflecting the nuanced requirements and challenges encountered at each stage of development. For example the C-RTL equivalence checking rigorously compares the RTL implementations against the C-based specification models. This verification method, as evidenced by studies such as [29, 30], is frequently applied, particularly in the context of data-path intensive designs, as highlighted by Hector from Synopsis[31]. Given that circuit designs within this abstract primarily encapsulate hardware behavior while abstracting concrete physical details, theorem provers and SMT solvers emerge as pivotal tools for enhancing verification efficacy [32, 33].
@@ 2.3 EDA for Back-End Design
For modern chip design, the back-end design flow, also referred to as layout design, is depicted in Fig. 2 b), transitioning from a gate-level or generic technology (GTech) netlist to a finalized layout [38].
@@ 2.4.1 EDA for Standard Cells
Standard cells, the building blocks of digital ICs, follow predefined structures that align with a library’s specifications, enabling their reuse across diverse designs. The focus of EDA tools in standard cell design is primarily on automating the layout generation process, encompassing crucial steps like placement and in-cell routing, as shown in Fig. 3.
The placement process is dedicated to determining the optimal transistor locations within a cell to maximize space utilization while maintaining functionality and performance integrity. The common solution algorithms for the placement includes dynamic program, reinforcement learning, and satisfiability modulo theories. Innovations in placement strategies, as highlighted in [40, 41], have introduced methods to expedite this intricate procedure while ensuring routability and design efficiency. In contrast, in-cell routing tackles the intricate task of establishing connections within the cell, a process complicated by the rigorous area constraints of standard cells. The in-cell routing are usually solved by A-star, interger linear programming, and satisfiability modulo theories. This stage demands specialized routing solutions, distinct from those applied to broader digital circuits, to navigate the tight confines of cell layouts. Contributions from [42, 43] have provided targeted approaches to in-cell routing, addressing the unique challenges of standard cell design.
@@ 2.4.2 EDA for Datapath Circuits
Figure 4: A typical datapath circuits design flow.
@@
These tools adopt a comprehensive strategy for refining datapath circuits, ensuring seamless integration and peak efficiency among components. As depicted in Fig. 4, the design journey initiates with pinpointing a target application and its corresponding architectural design, thereby defining a broad and intricate design space. This space might include, for example, CPU tasks like SPEC2017 benchmarks or GPU tasks such as matrix multiplication, targeting either CPU or GPU architectures accordingly. The design space diverges into two principal domains: the application space, outlining application-specific parameters like dataflow patterns or neural network mapping strategies, and the architecture space, detailing the structural and resource parameters, such as CPU issue width or the quantity of MACs in a neural processing unit (NPU).
The intersection of parameters from these domains establishes a “design point”, which, upon post-compilation application mapping, is subjected to thorough evaluation and validation via cutting-edge EDA tools. This rigorous process iteratively explores and assesses new design points until the exploration goals are achieved.
@@ 2.4.3 EDA for Analog Circuits
Fig. 5 illustrates a typical analog IC design flow, starting with a detailed set of circuit specifications covering area, power, and performance requirements. The front-end design phase is crucial, establishing the pre-layout circuit netlist that defines the circuit’s functionalities via meticulous topology design and device sizing. This phase sets the foundation for the circuit’s operational features and optimization criteria.
@@ 3.1.1 Pre-RTL ML Methods
In the realm of tensor computations, HASCO [78] employs ML for DSE. This methodology optimizes both software programs and hardware accelerators, showcasing ML’s capacity to bridge the gap between software and hardware domains to achieve optimized system performance.
@@ 3.1.2 RTL-Stage ML Methods
Power modeling at the RTL stage has also attracted lots of attention. There are two primary categories: design-time power estimation and runtime on-chip power modeling. For design-time estimation, PRIMAL [85] stands out for offering per-cycle power evaluations tailored to each target design, alongside other notable ML-based approaches [86, 87]. For runtime power modeling, DEEP [88] introduces an efficient on-chip model that incorporates low-overhead hardware design, utilizing ML to identify power-correlated RTL signals, or ‘power proxies’. This method, along with other ML-based on-chip power modeling solutions like [89, 90], demonstrates the potential of ML in creating dynamic power models that adapt to real-time conditions. Moreover, APOLLO [91] presents a versatile solution applicable to both design-time and runtime scenarios. Simmani [92] and the early power modeling work [93] focus on fast power emulation on FPGA and other platforms, highlighting the broader applicability of ML methods in facilitating efficient power analysis during the design phase.
@@ 3.1.4 Layout-Stage ML Methods
ML for Placement Stage Enhancements The placement stage, which determines the optimal locations of macros and standard cells in the layout, is pivotal for achieving desired design metrics. Early applications of ML aimed to augment traditional placement strategies. PADE [111] incorporates support vector machines (SVM) and neural networks for datapath extraction and evaluation, facilitating datapath-aware placement strategies. DREAMPlace, developed by Lin et al. [112], conceptualizes the placement challenge as akin to training a neural network, thus accelerating the global placement process by harnessing GPU computing capabilities. Building on DREAMPlace, Agnesina et al. [113] apply multi-objective Bayesian optimization for macro placement design space exploration, demonstrating the potential of ML in enhancing macro-placement outcomes.
@@
ML for Sign-Off Enhancements: During the routing and sign-off stages, the precision of sign-off timing, especially using the path-based static timing analysis (PBA), becomes crucial. However, the PBA process is time-consuming, leading to the application of machine learning models for predicting path-based timing based on quicker graph-based analysis (GBA) results. The pioneering work by Kahng et al. [129] was instrumental in predicting PBA from GBA using carefully engineered features and a tree-based model. Subsequent studies, such as  [124, 130], have delved into various machine learning models, including transformers and GNN, to enhance the accuracy of GBA-PBA predictions.
@@
ML for Manufacturability Enhancements: In the field of design for manufacturing (DFM), leveraging ML has become pivotal for bolstering the reliability of lithography and manufacturing processes, with layout patterns often analyzed as images. Studies like GAN-SRAF [134], GAN-OPC [135], Develset [136], and L2O-ILT [137] use various ML methods to improve mask synthesis printability. Other works, such as those by Watanabe et al. [138], Ye et al. [139], Lin et al. [140] and Chen et al. [141], focus on lithography modeling to simulate printed patterns from mask clips. For identifying layout patterns prone to printing failures like shorts or opens, ML-enhanced lithography hotspot detection is explored in various studies. For example, Yang et al. [142] propose to extract layout features with discrete cosine transform and utilize a CNN architecture for hotspot detection. The performance is further improved with the proposed bias learning algorithm because of the imbalanced dataset. Inspired by the object detection problem in computer vision, Chen et al. [143] propose to detect multiple hotspots within large layouts simultaneously. In [144], the binarized neural network is utilized to speed up the hotspot detection flow. New network architecture is designed based on residual networks to achieve higher detection accuracy and performance. Additionally, ML further contributes to yield estimation and analysis, as seen in works like Ciccazzo et al. [145], Nakata et al. [146], and Alawieh et al. [147].
@@ 3.2 Reinforcement Learning in EDA
Reinforcement learning (RL) in EDA has emerged as a powerful method for navigating the expansive solution spaces inherent in logic synthesis and physical design, often uncovering innovative solutions that surpass traditional, intuition-based approaches. Innovations like Synopsys.ai [152] underscore this trend, showcasing AI-driven methodologies that enhance PPA metrics across the design spectrum.
@@ 3.3 Leveraging Large Language Models in EDA
This section delves into the use of LLMs for RTL code generation—a key area of focus. It categorizes the research into LLM-aided RTL design generation and verification. Additionally, we explore LLM applications in generating EDA scripts and high-level architecture design.
@@ 3.3.2 Enhancing Verification with LLMs
Functional Verification through LLMs: LLMs have made significant strides in functional verification by translating natural language specifications into SystemVerilog assertions (SVAs). This process ensures that RTL implementations adhere to their intended specifications. Notably, [176, 177] leverage human-written specification sentences alongside RTL designs to generate precise SVAs. AssertLLM [178] takes a proactive approach by generating assertions directly from comprehensive specification documents, even before the RTL design phase. This method is complemented by a benchmark set that pairs natural language specifications with golden RTL implementations, offering a robust framework for evaluating assertion generation. Furthermore, LLMs have achieved success in solving the Boolean Satisfiability (SAT) problem [179], which can be applied to verify arithmetic circuits.